1) Calculate  the  heat  through  a  glass  window  that  is
30.0  cm*150  cm  in  area  and  1.20  mm  thick.  Assume  the
temperature onthe inside of the window is 25.0�C while the
outside temperature is 8.00�C.

2)Five hundred joules of heat is absorbed by a system that
does 200 J of work on its surroundings. What is the change in
the internal energy of the system?

  
3)An ideal gas expands isothermally, performing 8.80 kJ of
work  in  the  process.  Calculate  the  heat  absorbed  during  the
expansion.

4)A gas contained in a cylinder that has a piston is kept at
a  constant  pressure  of  2.80*10^5Pa. The  gas  expands  from0.500 m^3to 1.50 m^3
when 300 kJ of heat is added to the cylinder.
What is the change in internal energy of the gas?


5)The pressure in an ideal gas is slowly reduced to1/4
itsinitial value, while being kept in a container with rigid walls. In
the process, 800 kJ of heat leaves the gas. What is the change in
internal energy of the gas during this process?

6)Two  moles  of  an  ideal  monatomic  gas  expand  adiabati-
cally, performing 8.00 kJ of work in the process. What is the
change in temperature of the gas during the expansion?
Answer #1

1) dQ = k A dT/L = 1.05*.3*1.5*(25-8)/1.2E-3=6694

2) dU = Q- W = 500-200 = 300 J

3) isothermal so dU = 0
so Q = W
so Q = 8.8 kJ

4) dU = Q - W = 300.0E3 - 2.8E5*1= 20 kJ

5) since volume is constant W = 0
so dU = Q = -800 kJ

6)
adiabatic so Q = 0
dU = Q -W = -8.0 kJ